# Development and Characterization of MPGD-based Transition Radiation Detectors

### v1

@@ 4 Results
The primary performance parameters studied were: (i) gain stability under different high-voltage configurations, (ii) response to TR photons versus ionization along the track, (iii) timing characteristics of electron and pion signals between amplification technology choices, and (iv) e- separation quantified by suppression factors.
@@
At FTBF, systematic high-voltage scans were carried out for both the Micromegas-TRD and RWELL-TRD to map out operational stability. For the Micromegas-TRD, the drift field was scanned from about 1.3 to 1.5 kV/cm and the mesh voltage increased to a maximum of 675 V. Even at the maximum applied amplification voltage, the detector could not reach full efficiency and also suffered from a decrease in operational stability due to an increase in frequency of discharges. For the RWELL-TRD, the drift field was scanned up to 1.6 kV/cm, while the Micro-Well bias was increased to a maximum value of 540 V. Similar to the Micromegas-based prototype, operation at lower amplification voltage yielded stable but low-efficiency signals, whereas attempts to raise the voltage above 520 V produced frequent discharges without reaching an efficiency plateau.
@@
Despite visible TR photon clusters in both the Micromegas-TRD and RWELL-TRD, their gain was not sufficient to reach full efficiency and extract meaningful pion suppression factors. In contrast, the GEM-TRD_v1 exhibited stable operation across the full test campaign and served as the performance benchmark, achieving a pion detection efficiency measured to be 100% with an overall statistical uncertainty of 10%. Figure 7 (top) shows ADC spectra comparisons between the three prototypes at FTBF. Because the Micromegas prototype registered lower amplitudes than those of the GEM-TRD even at the highest amplification voltage application, a GEM preamplification layer was introduced prior to the CERN tests. At CERN, the Micromegas+GEM prototype showed improved gain and better agreement with the GEM-TRD_v2 reference technology (Figure 7 (bottom)). Both prototypes reached a detector efficiency plateau for all relevant charged particle species.
@@
Timing distributions were studied to distinguish TR clusters from the energy deposition along the track. Figure 8 shows the GEM-TRD_v1 drift-time response for 10 GeV electrons and pions with radiator material present. Ratios of radiator-to-no-radiator timing responses demonstrate clear excess electron signals, consistent with TR photon absorption predominantly near the cathode entrance - equivalent to energy deposits later in drift time. The timing response of the Micromegas+GEM-TRD for 20 GeV electrons and pions is shown in Figure 9. Compared to the GEM-TRD_v2, the overall signal extends over a longer drift time, consistent with the larger drift region of the Micromegas+GEM-TRD design. In addition, the leading and falling edge of the prototype’s distribution is less sharply defined than in the GEM-based prototype’s case. This behavior is expected, as the resistive amplification structure of a Micromegas mesh introduces slower charge collection relative to the fast induction signals that are characteristic of GEMs. The combination of extended drift length and a resistive amplification layer accounts for the broader timing profile observed in the Micromegas+GEM-TRD. In general, the addition of the GEM preamplification layer improved timing resolution and TR photon visibility relative to the Micromegas-TRD prototype’s performance at FTBF.
@@
Pion suppression factors were derived using neural network (NN) classifiers trained on amplitude and timing observables (ROOT-based TMVA root_tmlp). Ionization along each track was used as a neural network input layer, with the particle track drift time subdivided into several slices (sum of fADC samples). Cluster counting and characterization was also used as neural network input. A multilayer perceptron with two hidden layers was implemented for suppression factor determination. The data was split into two groups: one used for training, and another (independent sample) used for final decision evaluation. Figure 12, discussed in the next section, shows an example of the neural network output.
@@
At FTBF, the GEM-TRD_v1 achieved a pion suppression factor around 8 at 10 GeV for a representative working point of 90% electron efficiency with the fleece radiator (Fig. 10). It is important to note that this efficiency value is not a detector limitation, but rather one of several benchmark points used to quantify performance as depicted in Figure 10. No reliable suppression factors were extracted for the Micromegas-TRD or the RWELL-TRD at FTBF due to gain limitations previously discussed. At CERN, both the GEM-TRD_v2 and Micromegas+GEM-TRD were evaluated - Figure 11 shows the suppression factors for each detector obtained at 20 GeV for different radiator options. Note that the significant difference in each prototype’s performance with the foil-based radiator is explained by 1) the difference in overall radiator length and 2) the difference in construction as described in Section 3.3, with the foil-based radiator used on the Micromegas+GEM-TRD being better optimized for TR generation and transmission.
@@ 5 Discussion
Previous Monte Carlo simulations discussed in NIMGEMTRD1 predict a pion suppression factor of roughly 8 for a reference triple-GEM-TRD with 15 cm of radiator in Xe:CO2 90:10. The suppression performance of the GEM-TRD_v1 in 10 GeV mixed beam resulted in a suppression factor of about 8 with 25 cm of fleece radiator, while the CERN beam tests at 20 GeV yielded suppression factors of 4 for the GEM-TRD_v2 and 4.5 for the Micromegas+GEM-TRD at 90% electron efficiency with similar radiator settings. The authors attribute this significant discrepancy between previous results and these measurements primarily to the change in cathode material. Whereas the GEM-TRD_v1 prototype utilized a 0.2 m chromium cathode, both prototypes tested at CERN SPS employed a 5 m copper cathode. The TR photon spectrum relevant for absorption in Xe:CO2 spans roughly 3–40 keV, peaking near 11 keV. Copper has a strong K- absorption edge near 9 keV xray, meaning that the copper cathode absorbs a significant fraction of the TR energy spectrum exiting the radiator before it has a chance to enter the detector’s gas volume. This material choice would naturally reduce the number of detectable TR photons and therefore degrade the pion suppression factor. Chromium also has a K- absorption line near 6 keV, causing a similar though less significant effect in the GEM-TRD_v1 performance. Figure 12 contains an example of the NN output for the GEM-TRD_v2 based on electron and hadron signals collected at CERN SPS. The enhancement in the signal to the left of the 90% efficiency line is presumably a result of two separate occurrences: electrons which pass through the prototype without emission of TR photons, and electrons whose TR photon emission are not absorbed in the gaseous detector region.
@@
Monte Carlo studies using the Geant4 TR package geant4_trd were performed to model radiator photon yields and absorption in sequential layers of the detector prototypes. Figure 13 shows predicted TR photon spectra for 15 cm of regular-spaced radiator consisting of polypropylene foils of the same consistency as those used in the FTBF tests. Top plots in Figure 13 show results for the different GEM-TRD prototype constructions that were tested, for 10 GeV and 20 GeV electrons passing through, respectively. Bottom plots show results for 10 GeV and 20 GeV electrons passing through the Micromegas-TRD design and the Micromegas+GEM-TRD construction.
@@
These results emphasize that TRD optimization requires careful balancing of multiple design factors: radiator thickness and material for sufficient TR yield, cathode material composition and thickness for minimal soft TR x-ray stoppage, drift gap size for photon absorption depth, and amplification technology for signal timing and robustness. Future prototype iterations will need to co-optimize these parameters to achieve both efficient TR photon absorption and sharp particle separation capability. As a followup study, the authors have tested a variety of cathode materials on each prototype at Jefferson Lab in 3–6 GeV electron beam in order to directly demonstrate and quantify this effect. Analysis of these studies is currently underway and will be reported in forthcoming publications.
@@ 6 Summary and Future Work
We have reported on prototype development and beam test results for a triple-GEM-TRD, a Micromegas-TRD, a Micromegas+GEM-TRD, and a RWELL-TRD exposed to electron and pion beams at FTBF (10 GeV) and CERN SPS (20 GeV). The GEM-TRD_v1 achieved a pion suppression factor of about 8 at 90% electron efficiency with a 25 cm fleece radiator. Both the GEM-TRD_v2 and the Micromegas+GEM-TRD prototypes yielded significantly lower suppression factors - approximately 4 and 4.5, respectively, with a 23 cm fleece radiator. The principal factor identified is the cathode material, providing a natural explanation for the degraded suppression performance we observe. Prototypes tested in 20 GeV beam utilized a 5 m copper cathode, which eminently absorbs TR photons in the relevant energy range, thereby reducing the number of photons reaching the Xe:CO2 absorption region. Followup tests at Jefferson Lab have systematically varied cathode composition to quantify this impact and establish conditions under which suppression factors comparable to the GEM-TRD_v1 design can be determined.
Timing measurements further underscore the importance of amplification structure and drift configuration in a TRD application. The Micromegas-based prototype exhibited longer collection times and less sharply defined leading edges than the GEM-based prototypes, consistent with its larger drift gap and resistive amplification structure. These observations highlight an additional design trade-off: maximizing TR photon absorption through larger drift regions must be balanced against timing resolution for applications where charged particle tracking is a substantial component.
@@
The Micromegas-based TRD, with the addition of a GEM preamplification stage, showed clear improvement in stability and gain at CERN SPS relative to FTBF and demonstrated clear absorption and discrimination of TR photon signals to charged track dE/dx.
@@ Acknowledgments
The authors thank the staff at the Fermilab Test Beam Facility, including Mandy Rominsky, Eugene Schmidt, Joe Pastika, and Todd Nebel for their support and expertise during the beam test campaign. We extend similar thanks to the staff at the CERN SPS H8 beamline, including Barbara Holzer and Martin Jaekel. We also thank Matt Posik and Bernd Surrow of Temple University for technical contributions to the FTBF campaign, and the ATLAS TRT group, including Semen Doronin, Anatoli Romaniouk, Konstanin Vorobev, and Konstantin Zhukov for their expertise and involvement during the CERN SPS campaign. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177. This work was also supported in part by Department of Energy Award DE-FG05-92ER40712, DOE EIC Generic R&D 2022_02, and Vanderbilt University. Finally, the authors thank Beni Zihlmann and University of Virginia members Huong Nguyen and Nilanga Liyanage for their longtime support and involvement with these efforts.
Appendix A Triple-GEM-TRD Voltage Divider
| Equivalent divider current Ieq [A] | 343 | 354 | 362 |
| V across drift gap [V] | 3000 | 2900 | 3020 |
| V across GEM1 [V] | 409 | 426 | 439 |
| V across TG1 [V] | 694 | 711 | 724 |
| V across GEM2 [V] | 374 | 391 | 404 |
| V across TG2 [V] | 694 | 711 | 724 |
| V across GEM3 [V] | 338 | 355 | 368 |
| V across IG [V] | 694 | 711 | 724 |
| Drift gap E-field [kV/cm] | 1.43 | 1.38 | 1.52 |
| GEM1 E-field [kV/cm] | 81.8 | 85.2 | 87.8 |
| TG1 E-field [kV/cm] | 3.47 | 3.56 | 3.62 |
| GEM2 E-field [kV/cm] | 74.8 | 78.2 | 80.8 |
| TG2 E-field [kV/cm] | 3.47 | 3.56 | 3.62 |
| GEM3 E-field [kV/cm] | 67.6 | 71 | 73.6 |
| IG E-field [kV/cm] | 3.47 | 3.56 | 3.45 |
Appendix B Micromegas+GEM-TRD Voltage Divider
| Equivalent divider current Ieq [A] | 41 | 42 | 39.5 |
| V across drift gap [V] | 3470 | 3360 | 3420 |
| V across GEM1 [V] | 405 | 410 | 395 |
| V across TG1 [V] | 610 | 615 | 592 |
| Voltage on mesh [V] | 610 | 615 | 592 |
| Drift gap E-field [kV/cm] | 1.36 | 1.32 | 1.34 |
| GEM1 E-field [kV/cm] | 81.5 | 82 | 79 |
| TG1 E-field [kV/cm] | 2.45 | 2.46 | 2.37 |

### v2

@@ 4 Results
The primary performance parameters studied were: (i) operational stability under different high-voltage configurations, (ii) response to TR photons versus ionization along the track, (iii) timing characteristics of electron and pion signals between amplification technology choices, and (iv) e/ separation quantified by suppression factors.
@@
At FTBF, systematic high-voltage scans were carried out for both the Micromegas-TRD and RWELL-TRD to map out operational stability. For the Micromegas-TRD, the drift field was scanned from about 1.3 to 1.5 kV/cm and the mesh voltage increased to a maximum of 675 V. Even at the maximum applied amplification voltage, the maximum detector efficiency reached was below 80% and the detector also suffered from a decrease in operational stability due to an increase in frequency of discharges. For the RWELL-TRD, the drift field was scanned up to 1.6 kV/cm, while the Micro-Well bias was increased to a maximum value of 540 V. Similar to the Micromegas-based prototype, operation at lower amplification voltage yielded stable signals, whereas attempts to raise the voltage above 520 V produced frequent discharges with a detection efficiency below 70%.
@@
Despite visible TR photon clusters in both the Micromegas-TRD and RWELL-TRD, their signal amplitude was not sufficient to reach full efficiency and extract meaningful pion suppression factors. In contrast, the GEM-TRD_v1 exhibited stable operation across the full test campaign and served as the performance benchmark, achieving a pion detection efficiency measured to be 96.2% with an overall statistical uncertainty of 1.9%. Figure 7 (top) shows pulse amplitude spectra comparisons between the three prototypes at FTBF. Note that unlike Figure 4, which displays raw waveform sample distributions, Figure 7 shows differential spectra of pedestal-subtracted pulse peaks normalized to the number of selected events. Because the Micromegas prototype registered lower amplitudes than those of the GEM-TRD even at the highest amplification voltage application, a GEM preamplification layer was introduced prior to the CERN tests. At CERN, the Micromegas+GEM prototype showed improved gain and better agreement with the GEM-TRD_v2 reference technology (Figure 7 (bottom)). Both prototypes reached a detector efficiency plateau for all relevant charged particle species.
@@
Drift time distributions were studied to distinguish TR clusters from the energy deposition along the track. Figure 8 shows the GEM-TRD_v1 drift-time response for 10 GeV electrons and pions with radiator material present. Ratios of radiator-to-no-radiator signal shapes in time demonstrate clear excess electron signals, consistent with TR photon absorption predominantly near the cathode entrance — equivalent to energy deposits later in drift time. The drift time distributions of the Micromegas+GEM-TRD for 20 GeV electrons and pions are shown in Figure 9. Compared to the GEM-TRD_v2, the overall signal extends over a longer drift time, due in part to the larger drift region of the Micromegas+GEM-TRD design and the lower drift field strength. In addition, the leading and falling edge of the prototype’s distribution is less sharply defined than in the GEM-based prototype’s case; this behavior is expected, as the amplification structure of a Micromegas mesh introduces slower signal collection due to the ion charge contribution [11] relative to the fast induction electron signals that are characteristic of GEMs. In general, the addition of the GEM preamplification layer improved timing resolution and TR photon visibility relative to the Micromegas-TRD prototype’s performance at FTBF.
@@
Pion suppression factors were derived using neural network (NN) classifiers trained on amplitude and timing observables (ROOT-based TMVA [7]). Ionization along each track was used as a neural network input layer, with the particle track drift time subdivided into several slices (sum of fADC samples). Cluster counting and characterization was also used as neural network input. A multilayer perceptron with two hidden layers was implemented for suppression factor determination. The data was split into two groups: one used for training, and another (independent sample) used for final decision evaluation. Figure 12, discussed in the next section, shows an example of the neural network output.
@@
At FTBF, the GEM-TRD_v1 achieved a pion suppression factor around 8 at 10 GeV for a representative working point of 90% electron efficiency with the fleece radiator (Figure 10). It is important to note that this efficiency value is not a detector limitation, but rather one of several benchmark points used to quantify performance as depicted in Figure 10. No reliable suppression factors were extracted for the Micromegas-TRD or the RWELL-TRD at FTBF due to gain limitations previously discussed. At CERN, both the GEM-TRD_v2 and Micromegas+GEM-TRD were evaluated - Figure 11 shows the suppression factors for each detector obtained at 20 GeV for different radiator options. Note that the significant difference in each prototype’s performance with the foil-based radiator is explained by the difference in overall radiator length and construction as described in Section 3.2 and Table 2, with the foil-based radiator used on the Micromegas+GEM-TRD being better optimized for TR generation and transmission.
@@ 5 Discussion
Previous Monte Carlo simulations discussed in [4] predict a pion suppression factor of roughly 8 for a reference triple-GEM-TRD with 15 cm of radiator in Xe:CO2 90:10. The suppression performance of the GEM-TRD_v1 in 10 GeV mixed beam resulted in a suppression factor of about 8 with 25 cm of fleece radiator, while the CERN beam tests at 20 GeV yielded suppression factors of 4 for the GEM-TRD_v2 and 4.5 for the Micromegas+GEM-TRD at 90% electron efficiency with similar radiator settings. The authors attribute this significant discrepancy between previous results and these measurements primarily to the change in cathode material. Whereas the GEM-TRD_v1 prototype utilized a 0.2 m chromium cathode, both prototypes tested at CERN SPS employed a 5 m copper cathode. The TR photon spectrum relevant for absorption in Xe:CO2 spans roughly 3–40 keV, peaking near 11 keV. Copper has a strong K- absorption edge near 9 keV [8], meaning that the copper cathode absorbs a significant fraction of the TR energy spectrum exiting the radiator before it has a chance to enter the detector’s gas volume. This material choice would naturally reduce the number of detectable TR photons and therefore degrade the pion suppression factor. Chromium also has a K- absorption line near 6 keV, causing a similar though less significant effect in the GEM-TRD_v1 performance. Figure 12 contains an example of the NN output for the GEM-TRD_v2 based on electron and hadron signals collected at CERN SPS. The enhancement in the signal to the left of the 90% efficiency line is presumably a result of two separate occurrences: electrons which pass through the prototype without emission of TR photons, and electrons whose TR photon emission are not absorbed in the gaseous detector region.
@@
Monte Carlo studies using the Geant4 TR package [13] were performed to model radiator photon yields and absorption in sequential layers of the detector prototypes. Figure 13 shows predicted TR photon spectra for 15 cm of regular-spaced radiator consisting of polypropylene foils of the same consistency as those used in the FTBF tests. Top plots in Figure 13 show results for the different GEM-TRD prototype constructions that were tested, for 10 GeV and 20 GeV electrons passing through, respectively. Bottom plots show results for 10 GeV and 20 GeV electrons passing through the Micromegas-TRD design and the Micromegas+GEM-TRD construction.
@@
These results emphasize that TRD optimization requires careful balancing of multiple design factors: radiator thickness and material for sufficient TR yield, cathode material composition and thickness for minimal soft TR X-ray stoppage, drift gap size for photon absorption depth, and amplification technology for signal timing and robustness. This work stands to establish clear design requirements for MPGD-based TRD operation and demonstrates that multi-stage amplification is not merely advantageous but necessary for achieving meaningful TRD performance with MPGDs. As a followup study, the authors have tested a variety of cathode materials on each prototype at Jefferson Lab in 3–6 GeV electron beam in order to directly demonstrate and quantify this effect. This ongoing R&D effort will also focus on refining these multi-stage-gain detector configurations to maximize suppression performance while minimizing material and complexity, guiding the development of next-generation TRDs for high-rate experimental environments. Analysis of these studies is currently underway and will be reported in forthcoming publications.
@@ 6 Summary and Future Work
We have reported on prototype development and beam test results for two triple-GEM-TRDs, a Micromegas-TRD, a Micromegas+GEM-TRD, and a RWELL-TRD exposed to electron and pion beams at FTBF (10 GeV) and CERN SPS (20 GeV). The GEM-TRD_v1 achieved a pion suppression factor of about 8 at 90% electron efficiency with a 25 cm fleece radiator. Both the GEM-TRD_v2 and the Micromegas+GEM-TRD prototypes yielded significantly lower suppression factors - approximately 4 and 4.5, respectively, with a 23 cm fleece radiator. The principal factor identified is the cathode material, providing a natural explanation for the degraded suppression performance observed. Prototypes tested in 20 GeV beam utilized a 5 m copper cathode, which eminently absorbs TR photons in the relevant energy range, thereby reducing the number of photons reaching the Xe:CO2 absorption region. Followup tests at Jefferson Lab have systematically varied cathode composition to quantify this impact and establish conditions under which suppression factors comparable to the GEM-TRD_v1 design can be determined.
Drift time measurements further underscore the importance of amplification structure and drift configuration in a TRD application. The Micromegas-based prototype exhibited longer collection times and less sharply defined leading edges than the GEM-based prototypes, consistent with its larger drift gap and slower ion-dominated signal formation. These observations highlight an additional design trade-off: maximizing TR photon absorption through larger drift regions must be balanced against timing resolution for applications where charged particle tracking is a substantial component.
@@
The Micromegas-based TRD, with the addition of a GEM preamplification stage, showed clear improvement in operational stability and gain at CERN SPS relative to FTBF and demonstrated clear absorption and discrimination of TR photon signals to charged track dE/dx.
@@ Acknowledgments
The authors thank the staff at the Fermilab Test Beam Facility, including Mandy Rominsky, Eugene Schmidt, Joe Pastika, and Todd Nebel for their support and expertise during the beam test campaign. We extend similar thanks to the staff at the CERN SPS H8 beamline, including Barbara Holzer and Martin Jaekel. We also thank Matt Posik and Bernd Surrow of Temple University for technical contributions to the FTBF campaign, and the ATLAS TRT group, including Semen Doronin, Anatoli Romaniouk, Konstanin Vorobev, and Konstantin Zhukov for their expertise and involvement during the CERN SPS campaign. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177. This work was also supported in part by Department of Energy Award DE-FG05-92ER40712, DOE EIC Generic R&D 2022_02, and Vanderbilt University. The work of A.A. was supported by the DOE, Office of Science, Office of Nuclear Physics in the Early Career Program. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program administered by the Oak Ridge Institute for Science and Education for the DOE under contract number DE‐SC0014664. Finally, the authors thank Beni Zihlmann from Jefferson Lab and University of Virginia members Huong Nguyen and Nilanga Liyanage for their longtime support and involvement with these efforts.
Appendix A TR Radiators
See Table 2.
| GEM-TRD_v1 | GEM-TRD_v2 | Micromegas-TRD | Micromegas+GEM-TRD | |||||
| Parameter | ||||||||
| Beam energy | 10 GeV | 10 GeV | 20 GeV | 20 GeV | 10 GeV | 10 GeV | 20 GeV | 20 GeV |
| Rad. material | Mylar | Fleece | Mylar | Fleece | Mylar | Fleece | Mylar | Fleece |
| Rad. layer thickness [m] | 12.5 | 2500 | 30 | 2500 | 12.5 | 2500 | 30&25 | 2500 |
| Gap material | Netted spacer | Air | Air | Air | Air | Netted spacer | Air | Air |
| Gap thickness [m] | 300 | N/A | 240 | N/A | 300 | N/A | 200&200 | N/A |
| N layers | 474 | 100 | 555 | 60, 90 | 474 | 100 | 450&450 | 60, 90 |
| Overall length [cm] | 25 | 25 | 15 | 15, 23 | 25 | 25 | 21 | 15, 23 |
Appendix B High Voltage Field Configurations
| Xe:CO2 Settings | |||
| Quantity | GEM-TRD_v1 | GEM-TRD_v2 | |
| Drift gap E-field [kV/cm] | 1.38 | 1.52 | |
| GEM1 E-field [kV/cm] | 85.2 | 87.8 | |
| TG1 E-field [kV/cm] | 3.56 | 3.62 | |
| GEM2 E-field [kV/cm] | 78.2 | 80.8 | |
| TG2 E-field [kV/cm] | 3.56 | 3.62 | |
| GEM3 E-field [kV/cm] | 71 | 73.6 | |
| IG E-field [kV/cm] | 3.56 | 3.45 | |
| Effective Gain | |||
| Xe:CO2 Settings | ||
| Quantity | ||
| Drift gap E-field [kV/cm] | 1.36 | |
| GEM E-field [kV/cm] | 81.5 | |
| TG E-field [kV/cm] | 2.45 | |
| Voltage on mesh [V] | 610 | |
| Effective Gain | ||